OpenStudy (anonymous):

Evaluate and leave the solution in standard notation: (2 x 10^4)(5.1 x 10^2) A. 10.2 B. 10,200,000 C. 1,200,000 D. 102,000

OpenStudy (anonymous):

= 10.2 * 10^6 = 1 .02 * 10^7 = 1,200,00 B

OpenStudy (anonymous):

thats 1,200,000 B